OpenStudy (anonymous):

how do you graph: Eq. 1: 2x - y= 8 Eq. 2: 3x - 7= 4y

OpenStudy (anonymous):

Find x and y intercepts for each line and graph each line in two dimensional space on x/y axes.

OpenStudy (anonymous):

Do you know how to find x and y intercepts?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Good. Can you work it and show me what you got?

OpenStudy (anonymous):

y=mx +b 2x - y = 8 y= 2x + -8 slope: 2 yint: -8 x int: 4 3x - 7= 4y slope: 3/4 im stuck on the yint and xint on 3x - 7= 4y

OpenStudy (anonymous):

for the y intercept let x = 0 and solve for y

OpenStudy (anonymous):

you get 0 - 7 = 4y

OpenStudy (anonymous):

so y = -7/4

OpenStudy (anonymous):

Could you get by yourself the x intercept?

OpenStudy (anonymous):

i think you can by multiplying both sides by 3/4

OpenStudy (anonymous):

Multiply both sides of the second equation you mean.

OpenStudy (anonymous):

And what is the next step?

OpenStudy (anonymous):

umm you cross cancel?